8/31/16

Summary:

No winds this morning. Notice no pressure coming in from s3 and s15. The xml config needs to be updated.

Actions past 24 hours:

To dos:

Sensor status:

T/RH:  s9,s10 down
P: ok at s1
2D Gill: ok
csat u,v:  ok
csat ldiag: spiking up to 1
csat w, tc: ok
EC150: ok
motes: data not coming in
Wetness: unknown
radiation: unknown
Tsoil: unknown
Gsoil: unknown
Qsoil: unknown
Cvsoil: unknown
Battery: Gary says he needs to do some configuration work